OpenStudy (anonymous):

using the properties of exponents and radicals, design some different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Pleaseeee help me on how to do this

OpenStudy (anonymous):

x * x = x^1 * x^1 = x^2 |sqrt (x^4)| = +x^(4/2) = x^2 x^6/x^4 = x^(6-4) = x^2 on and on and on :)

OpenStudy (anonymous):

thank you!